Use the​ Gauss-Jordan method to solve the following system of equations.

5x+4y-3z+0

2x-y+5z=1

7x+3y+2z=1

Multiple Choice

A.The solution is

B.There is an infinite number of solutions. The solution is

C.

There is no solution.
Answer #1

5x + 4y - 3z = 0 - (1) 2x - y + 5z = 1 -(2) 7x + 3y – 2z = 1 - (3) Converting given equations into matrix form 54 - 30 2 -1 54 R; --21-5*R R:=R 0 0 0 0 17 x132 4 13 4 17 *13 13- 5 31 y = - 13 13 Solution By Gauss jordan elimination method 4 17 5 31 *Since we get solution in parametric terms of z.

So, we have infinite many solution. Option B is correct.

Please give it a thumbs up if you like the answer. Comment if any problem in solution. :)